OpenStudy (anonymous):

y = f(x) = 2^x Find f(x) when x = 1

OpenStudy (aum):

Just put x = 1 \[ f(x) = 2^x \\ f(1) = 2^1 = 2 \]

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (aum):

yes.

OpenStudy (anonymous):

thank you !